OpenStudy (anonymous):

What is the value of x in the solution to the following system of equation x - y = -3 x + 3y = 5

OpenStudy (anonymous):

from first equation you will get y=x+3 put that value in 2nd equation and find value of x

OpenStudy (anonymous):

@aajugdar so it would be x + y = 2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

@maddie3 Oh thank uu. I appreciate that!

OpenStudy (anonymous):

x+ 3(x+3) =5 x+3x+9 =5 4x=5-9 4x=-4 x=-1